# From Trace to Line: LLM Agent for Real-World OSS Vulnerability Localization

### v2

@@ Abstract
Large language models show promise for vulnerability discovery, yet prevailing methods inspect code in isolation, struggle with long contexts, and focus on coarse function or file level detections which offers limited actionable guidance to engineers who need precise line-level localization and targeted patches in real-world software development. We present T2L-Agent (Trace-to-Line Agent), a project-level, end-to-end framework that plans its own analysis and progressively narrows scope from modules to exact vulnerable lines. T2L-Agent couples multi-round feedback with an Agentic Trace Analyzer (ATA) that fuses runtime evidence such as crash points, stack traces, and coverage deltas with AST-based code chunking, enabling iterative refinement beyond single pass predictions and translating symptoms into actionable, line-level diagnoses. To benchmark line-level vulnerability discovery, we introduce T2L-ARVO, a diverse, expert-verified 50-case benchmark spanning five crash families and real-world projects. T2L-ARVO is specifically designed to support both coarse-grained detection and fine-grained localization, enabling rigorous evaluation of systems that aim to move beyond file-level predictions. On T2L-ARVO, T2L-Agent achieves up to 58.0% detection and 54.8% line-level localization, substantially outperforming baselines. Together, the framework and benchmark push LLM-based vulnerability detection from coarse identification toward deployable, robust, precision diagnostics that reduce noise and accelerate patching in open-source software workflows. Our framework and benchmark are publicly available as open source at https://github.com/haoranxi/T2LAgent.
@@ 1 Introduction
Software vulnerabilities now occur at unprecedented scale and cost. In 2023, more than 29,000 Common Vulnerabilities and Exposures (CVEs) were recorded cve_details_2024. In the first half of 2025, 1,732 data breaches were reported–an 11% increase indusface_2024. The economic toll is mounting: software supply-chain attacks are projected to cost the global economy 80.6 billion annually by 2026, up from 45.8 billion in 2023 darkreading_supply_chain; some estimates put total damages at \$9.5 trillion in 2024 liu_makridis_galinkin_2024. Critically, 14% of breaches in 2024 began with vulnerability exploitation—nearly triple the prior year. Despite advances in automated vulnerability detection, effectiveness in localization remains underexplored zhang2024empirical, leaving developers with coarse, file-level predictions while 32% of critical vulnerabilities remain unpatched for over 180 days. These vulnerabilities are exploding in volume and growing ever more severe, driving escalating security and economic risks worldwide.
@@
Motivation. The rise of Large Language Models (LLMs) has accelerated AI-driven automation across software development and cybersecurity zhang_et_al_2025. From code assistants like github_copilot_2021 to automated security analysis and automation nunez2024autosafecoder. LLMs show strong aptitude for understanding complex codebases and flagging potential issues divakaran2024llms. This momentum naturally extends to vulnerability detection, spanning model-level approaches such as fine-tuning LLMs for classification—and agentic frameworks that harness LLM reasoning for automated analysis awesome_llm4cybersecurity_2024.
Yet today’s LLM-based detection methods face practical barriers. Most operate at the function level, asking whether a fragment is vulnerable rather than pinpointing where the flaw lies zhang2024empirical, 10.1145/3715758. These tasks are also defined purely on static code, ignoring binaries, runtime behavior, and project-level diffs, further limiting their usefulness for real remediation. Evaluations also rely heavily on lightweight, synthetic datasets that miss the complexity of production systems springer_vulnerability_analysis.
This research–practice gap is most visible in vulnerability localization. While studies show promising results on isolated functions or snippets, engineers must navigate large repositories with cross-module dependencies and need line-level localization to craft minimal, targeted patches precision current methods rarely deliver. Moreover, artificial benchmarks obscure the true difficulty of project-level detection, where context spans multiple files and demands system-wide reasoning.
Given the surge in vulnerability reports outlined earlier, the field urgently needs approaches that close this reality gap: leveraging LLM automation while tackling project-scale challenges practitioners face daily. Only by embracing these real-world constraints can we convert the promise of LLM-assisted security into tools that meaningfully lighten the load on development teams.
@@
To address these gaps, we introduce T2L (Trace-to-Line), which reframes vulnerability detection as a two-tier problem (a) coarse-grained detection - flagging suspicious code chunks and (b) fine-grained localization - pinpointing exact vulnerable lines. This separation enables systematic evaluation of LLM capabilities from repository-scale reasoning to human-expert precision.
@@
Contribution. This work makes three contributions: (1) New Task formulation. We define a new task for agentic AI, a runtime trace guided, project-level vulnerability detection, and formulate it as two structured subtasks: chunk-level detection and line-level localization. AST-based chunking adapts large codebases to LLM context while preserving semantics, enabling refinement from coarse predictions to exact lines, bridging research setups and real-world needs for LLMs. (2) T2L-ARVO benchmark. We present the first benchmark for agentic fine-grained localization, featuring 50 expert verified cases across five vulnerability types with balanced category distribution. T2L-ARVO enables realistic, project-scale evaluation of LLM-based systems. (3) T2L-Agent framework. We propose a multi-agent system with the following innovation: (i) a Trace Analyzer (Agentic Trace Analyzer) integrating the ensemble of tools such as static analysis, sanitizers, and runtime monitoring for observability as human debuggers; (ii) a Proposal module (Divergence Tracing) that iteratively forms and tests vulnerability hypotheses; and (iii) a two-stage pipeline (Detection Refinement) that first detects coarse chunks and then verifies exact lines with runtime feedback.
Overall, T2L-Agent provides a comprehensive baseline for this new vulnerability localization task. Unlike prior snippet-level detectors, it is designed to operate on fully reproducible real-world projects, using crashes, sanitizer reports, and project diffs as runtime evidence. On T2L-ARVO, it achieves up to 58% chunk-level detection and 54.8% exact line localization, showing that this runtime-evidence–guided setting is challenging yet tractable for LLM agents.
@@ 2 Background
Vulnerability Localization. Classical localization combines static and dynamic analyses with retrieval and graph methods. Static slicing selects statements that may affect a slicing criterion to form behavior-preserving slices for debugging and comprehension 10.5555/800078.802557. Dynamic slicing refines this by deriving input-specific slices from execution-time dependence graphs, yielding higher precision near crashes or variables agrawal1990dynamic. Information-retrieval approaches treat bug reports as queries, ranking files/methods by textual and historical signals to surface likely locations zhou2012should; 6693093; 10.1145/2597008.2597148. Code Property Graphs unify classic program-analysis concepts in a joint representation, enabling scalable pattern searches that uncovered 18 previously unknown vulnerabilities yamaguchi2014modeling.
AI for Cybersecurity. AI is widely applied to cybersecurity tasks with tool execution and feedback in runnable environments yang2023intercode. One prominent use is offensive security shao2024empirical, where agents reason about attack chains and interact with benchmark environments shao2024nyu; zhang2024cybench to simulate human CTF players udeshi2025d; shao2025craken; abramovichenigma. In parallel, pen-testing systems orchestrate modules or multi-agent collaboration to automate reconnaissance, vulnerability analysis, and exploitation against targets deng2024pentestgpt; shen2025pentestagent. Further work builds agent-based patching frameworks that localize vulnerabilities and synthesize fixes yu2025patchagent; xue2025pagent, and explores LLM-augmented red and blue teaming, where agents emulate adaptive attackers and assist defenders in threat hunting and incident investigation abuadbba2025promise; liu2025benchmarking.
LLM Agentic Systems. Beyond cybersecurity, large language models (LLMs) are increasingly applied to automate workflows across diverse domains shao2024survey, including AI for scientific research bran2023chemcrow; basit2025pennylang; jin2024prollm, software engineering yang2024swe, healthcare isern2016systematic, and hardware design wang2024llms. Many systems integrate long-term memory to retain intermediate results, environment states, and user preferences across sessions zhang2025survey. LLM as judge are often used to evaluate actions, select among tool outputs, and provide feedback for self-improvement  li2025generation. Some agents employ self-reflection and critique loops to iteratively refine their reasoning or code before execution renze2024self. Multi-agent and role-specialized architectures further enable collaboration among agents with distinct skills to solve complex, interleaved subtasks in parallel li2024survey.
Figure 1: T2L-Agent Framework overview.
@@ 3 Related Work
LLMs show strong potential on vulnerability localization tasks. Early learning-based detectors improve localization by training models to classify whether a unit is vulnerable. LLMAOyang2024large fine-tunes LLMs on small, manually curated buggy programs, while BAPstein2025s learns state-of-the-art vulnerability localization without any direct localization labels, outperforming traditional baseline over eight benchmarks. However, subsequent analyses revealed data issues in widely used datasets such as Big-Vul and Devignzhou2019devign, making people doubt about performance numbers and highlighting the need for realistic evaluation settings croft2023data.
Study
Line Lv.
Mult. Ag.
Runtime
Iterative
LLMAO yang2024large
BAP stein2025s
GenLoc asad2025leveraging
AgentFL qin2024agentfl
CoSIL jiang2025cosil
AutoFL kang2024quantitative
LineVul 9796256
LOVA li2024attention
MatsVD 10.1145/3671016.3674807
xLoc 10.1145/3660804
LLM4FL rafi2024multi
MemFL yeo2025improving
@@
Figure 2: Related Works
Many prior works frame vulnerability localization at the file or function level, but this coarse granularity often fails to offers limited guidance for developers. GenLoc asad2025leveraging identifies potentially vulnerable files from bug reports and iteratively analyzes them using code exploration tools. AgentFL qin2024agentfl applies a multi-agent framework for function-level localization, modeling the task as a three-step pipeline with specialized agents and tools. CoSIL jiang2025cosil narrows the function-level search space using module call graphs and iteratively traverses them for relevant context. Similarly, AutoFL kang2024quantitative prompts LLMs to localize method-level vulnerabilities via function-call navigation, showing that multi-step reasoning mitigates context limits.
Recent studies have shifted toward line or statement level localization. LineVul 9796256 uses a Transformer-based classifier for line-level prediction, while LOVA li2024attention introduces a self-attention framework to score and highlight vulnerable lines. MatsVD 10.1145/3671016.3674807 enhances statement level localization with dependency-aware attention, and xLoc 10.1145/3660804 learns multilingual, task-specific knowledge for bug detection and localization. Building on these efforts, LLM4FL rafi2024multi proposes a multi-agent framework leveraging graph-based retrieval and navigation to reason about failure causes. MemFL yeo2025improving introduces external memory to incorporate project-specific knowledge, improving localization in complex, repository-scale systems. CThey push localization to line level and use more multi-agent coordination. However, most still rely on limited runtime evidence, single-pass predictions, or benchmarks that lack realistic project settings.
T2L formalizes a new task of runtime-trace-guided vulnerability localization at the project level. Prior learning-based approaches such as BAP, LLMAO, and LineVul rely on static code analysis, training models to classify whether a given code snippet is vulnerable based solely on syntactic and semantic patterns. These methods typically operate on isolated single files and produce coarse-grained predictions at the function or file level without leveraging dynamic execution information. In contrast, our task targets multi-file codebases and incorporates runtime evidence including reproducible crash logs, stack traces, and sanitizer reports to pinpoint exact vulnerable lines. This formulation more closely reflects real-world vulnerability triage workflows where developers rely on both static code inspection and dynamic program behavior to diagnose security flaws.
@@ 4.1 T2L Agent
T2L-Agent uses a hierarchical planner–executor design udeshi2025d that splits localization into evidence gathering, hypothesis generation, and iterative refinement.Unlike single-pass analyzers, it uses a human-like process: gather runtime signals, link them to code, and narrow the search space. Details of the tool list that T2L originally supports are provided in  A.5.
Figure 3: ATA Components
Evidence Tracing T2L Planner coordinates repository analysis and runtime evidence capture in a single, structured pipeline. First, code-structure analysis partitions the codebase into function-aligned, semantically coherent units that preserve syntactic relationships while fitting LLM context windows; known patch locations are also indexed to establish evaluation baselines. Next, runtime evidence becomes the cornerstone: Sanitizer records memory-violation patterns, allocation traces, and stack frames, while interactive debugging (GDB/LLDB) provides symbolic context via backtraces and variable-state snapshots at crash points. This dual-layer design yields comprehensive, observable crash logs rather than speculative static signals. We collect multiple types of runtime and static evidence, including crash logs, sanitizer reports, stack traces, and static analyzer results, all evidence is merged into a single unified textual block and fed to the LLM at once. This design ensures that the LLM forms its reasoning based on the global evidence context, rather than aggregating per-evidence predictions to avoids issues of cross-signal inconsistency. The integrated workflow is a key T2L innovation; we detail it in Sec. 3.
Hypothesis Generation A “hypothesis” an LLM-generated candidate vulnerability produced from the unified evidence block of an existing vulnerable project in T2L system. The LLM proposes a list of file:line candidate pairs ranked by confidence. Each hypothesis is independently checked against the ground-truth patch diff, and candidates do not affect one another during evaluation.
Two-stage Refinement To ensure accurate detection, T2L uses a two-stage refinement process. It begins with coarse-grained filtering: the system derives broad candidate regions from crash logs, using LLM code comprehension to link symptoms to likely causes and produce a ranked list of candidates with confidence and rationale. It then identifies code fragments the LLM flagged as relevant in this first pass–such as missing bounds checks, uninitialized variables, or improper memory management. Next comes fine-grained refinement: these extracted fragments are added to the evidence for a second LLM pass, which updates the candidate set. Any newly generated locations are merged into the existing candidates. This refinement loop continues until the LLM produces no additional candidates or the cost budget is reached.
Feedback Control Each cycle outputs a brief task summary with success indicators and confidence. The Planner adapts the next step—continue refining or stop, preventing premature termination and avoiding over-analysis. The loop improves precision while controlling compute cost.
@@
Agentic Trace Analyzer (ATA) The Agentic Trace Analyzer (ATA) is the core module of T2L-Agent. Its purpose is to connect runtime behavior with relevant source code regions, addressing a gap in prior work that relies mainly on static signals. ATA constructs a unified “evidence block” by running targets in reproducible Docker, partitioning repositories into semantically meaningful chunks using tree-sitter, and instrumenting executions with analysis toolkit such as Sanitizers, Debuggers, and Static analyzers to collect stack traces, memory errors, and other runtime cues. The output is a unified “evidence block” which is then passed to T2L-Agent’s reasoning modules.
We use hierarchical refinement, starting from crash logs and narrowing to specific code. The agent cross-validates runtime signals with static features and ranks candidates by LLM confidence. Like human debugging, T2L-Agent seeds candidates via static–dynamic correlation and iteratively refines them with source–feedback loops. ATA brings fewer single-shot failures, improved compute efficiency, and behavior-anchored decisions that enable precise, line-level localization even in large, tightly coupled codebases. With ATA disabled, the LLM localizes on its own, as shown in Figure 4(c). Without ATA, the LLM fails to localize vulnerabilities compared with (a) and (b).
(a) Detection Refinement
(b) Divergence Tracing
@@
Figure 4: Partial T2L-Agent logs to show the how the three proposed technique on T2L-Agent work and help the task: Detection Refinement, Divergence Tracing and Agentic Trace Analyzer.

### v3

@@ Abstract
Large language models show promise for vulnerability discovery, yet prevailing methods inspect code in isolation, struggle with long contexts, and focus on coarse function- or file-level detections that offer limited guidance to engineers who need precise line-level localization for targeted patches. We introduce T2L, an executable framework for project-level, line-level vulnerability localization that progressively narrows scope from repository modules to exact vulnerable lines via AST-based chunking and evidence-guided refinement. We provide a baseline agent with an Agentic Trace Analyzer (ATA) that fuses runtime evidence such as crash points and stack traces to translate failure symptoms into actionable diagnoses. To enable rigorous evaluation, we introduce T2L-ARVO, an expert-verified 50-case benchmark spanning five crash families in real-world projects. On T2L-ARVO, our baseline achieves up to 58.0% detection and 54.8% line-level localization rate. Together, T2L framework advance LLM-based vulnerability detection toward deployable, precision diagnostics in open-source software workflows. Our framework and benchmark are publicly available as open source at https://github.com/haoranxi/T2LAgent.
Machine Learning, ICML
@@ 1 Introduction
Software vulnerabilities are widespread and costly, yet automated methods often provide coarse, file- or function-level predictions that are hard to turn into targeted fixes. Large language models (LLMs) have shown promise for code comprehension and vulnerability discovery, but today’s LLM-based approaches still face practical barriers for real remediation. In 2023, more than 29,000 CVEs were recorded (CVE Details, 2024). The economic toll is mounting: software supply chain attacks are projected to cost the global economy 80.6 billion annually by 2026 (Dark Reading, 2024). Critically, 14% of breaches in 2024 began with vulnerability exploitation, nearly triple the prior year. Despite advances in automated detection, localization remains underexplored (Zhang et al., 2024b), leaving developers with coarse predictions while 32% of critical vulnerabilities remain unpatched for over 180 days.
@@
Motivation. The rise of LLMs has accelerated AI-driven automation across software development and cybersecurity (Zhang et al., 2025a), from code assistants (GitHub, Inc., 2021) to automated security analysis (Nunez et al., 2024). LLMs show strong aptitude for understanding complex codebases and flagging potential issues (Divakaran and Peddinti, 2024), naturally extending to vulnerability detection through fine-tuning and agentic frameworks (tmylla and contributors, 2024).
Yet today’s LLM-based methods face practical barriers. Most operate at the function level, asking whether a fragment is vulnerable rather than pinpointing where the flaw lies (Zhang et al., 2024b; Sovrano et al., 2025). These tasks are defined mainly on static code, ignoring runtime behavior and project-level context. Evaluations rely on synthetic datasets that miss the complexity of production systems (Guo et al., 2024), while real-world debugging requires project-scale reasoning across multiple files. Engineers must navigate large repositories with cross-module dependencies and need line-level localization to craft targeted patches. Moreover, vulnerability analysis is often posed as static classification, ignoring diagnostic evidence such as crashes, sanitizer reports, and stack traces. As a result, the research-practice gap is most visible in localization: engineers need line-level guidance to craft patches, yet current systems rarely provide precise, actionable locations in complex repositories.
@@
To address these gaps, we introduce T2L (Trace-to-Line), which reframes vulnerability localization as a two-tier problem: (a) coarse-grained detection: flagging suspicious code chunks, and (b) fine-grained localization: pinpointing exact vulnerable lines. This separation enables systematic evaluation of LLM capabilities from repository-scale narrowing to expert-level line pinpointing.
@@
Contribution. This work makes three contributions: (1) New task formulation. We define runtime-trace-guided, project-scale vulnerability localization as two structured subtasks: chunk-level detection and line-level localization, enabling evidence-guided refinement from coarse predictions to exact lines. (2) T2L-ARVO benchmark. We present the first benchmark for agentic fine-grained localization, featuring 50 expert-verified cases across five vulnerability types with balanced category distribution. T2L-ARVO enables realistic, project-scale evaluation of LLM-based systems. (3) Evaluation framework and agent baseline. We introduce an evaluation framework that standardizes environment setup, evidence collection, and line-level metrics, together with an agent baseline integrating an Agentic Trace Analyzer, iterative hypothesis generation, and two-stage refinement. On T2L-ARVO, our baseline achieves up to 58% chunk-level detection and 54.8% exact line localization.
@@ 2 Background
Vulnerability Localization. Classical localization combines static and dynamic analyses with retrieval and graph methods. Static slicing selects statements affecting a slicing criterion for debugging (Weiser, 1981). Dynamic slicing derives input-specific slices from execution-time dependence graphs (Agrawal and Horgan, 1990). Information-retrieval approaches rank files/methods by textual and historical signals (Zhou et al., 2012; Saha et al., 2013; Wang and Lo, 2014). Code Property Graphs unify program-analysis concepts for scalable pattern searches (Yamaguchi et al., 2014).
AI for Cybersecurity. AI is widely applied to cybersecurity with tool execution in runnable environments (Yang et al., 2023). Offensive security agents (Shao et al., 2024b) reason about attack chains and interact with benchmarks (Shao et al., 2024c; Zhang et al., 2024a) to simulate CTF players (Udeshi et al., 2025; Shao et al., 2025; Abramovich et al., 2025). Pen-testing systems automate reconnaissance and exploitation (Deng et al., 2024; Shen et al., 2025). Agent-based frameworks localize vulnerabilities and synthesize fixes (Yu et al., 2025; Xue et al., 2025), while LLM-augmented red/blue teaming assists threat hunting (Abuadbba et al., 2025; Liu et al., 2025).
LLM Agentic Systems. Beyond cybersecurity, LLMs automate workflows across diverse domains (Shao et al., 2024a), including scientific research (Bran et al., 2023; Basit et al., 2025; Jin et al., 2024), software engineering (Yang et al., 2024c), healthcare (Isern and Moreno, 2016), and hardware design (Wang et al., 2024). Systems integrate memory for retaining states across sessions (Zhang et al., 2025b). LLM-as-judge evaluates actions and provides feedback (Li et al., 2025). Agents employ self-reflection to refine reasoning (Renze and Guven, 2024), and multi-agent architectures enable collaboration on complex subtasks (Li et al., 2024a).
Figure 1: T2L-Agent Framework overview. Left: input from T2L ARVO dataset. Center: Agentic Trace Analyzer (ATA) with Planner and Executor agents. Right: running demo showing the iterative planning-execution workflow for vulnerability localization.
@@ 3 Related Work
LLMs show strong potential for vulnerability localization. LLMAO (Yang et al., 2024a) fine-tunes LLMs on manually curated buggy programs, while BAP (Stein et al., 2025) learns localization without direct labels, outperforming baselines over eight benchmarks. However, data issues in Big-Vul and Devign (Zhou et al., 2019) raise concerns about evaluation validity (Croft et al., 2023).
Many works frame localization at file or function level, offering limited guidance for developers. GenLoc (Asad et al., 2025) identifies vulnerable files from bug reports iteratively. AgentFL (Qin et al., 2024) applies multi-agent frameworks for function-level localization. CoSIL (Jiang et al., 2025) narrows search space using module call graphs. AutoFL (Kang et al., 2024) prompts LLMs via function-call navigation, showing multi-step reasoning mitigates context limits.
Recent studies shift toward line-level localization. LineVul (Fu and Tantithamthavorn, 2022) uses Transformer-based classifiers, LOVA (Li et al., 2024b) introduces self-attention scoring, MatsVD (Weng et al., 2024) adds dependency-aware attention, and xLoc (Yang et al., 2024b) learns multilingual knowledge. LLM4FL (Rafi et al., 2024) leverages graph-based retrieval, and MemFL (Yeo et al., 2025) incorporates external memory for repository-scale systems. However, most still rely on limited runtime evidence or benchmarks lacking realistic project settings.
T2L formalizes trace-guided, project-scale, line-level localization. Unlike LLMAO, BAP, and LineVul that rely on static signals and operate on single files, our task targets training-free pipelines on multi-file codebases, incorporating runtime evidence including crash logs, stack traces, and sanitizer reports to pinpoint exact vulnerable lines. This formulation reflects real-world triage workflows where developers rely on both static inspection and dynamic behavior.
Table 1: Related Works
LLMAO 2024a
BAP 2025
GenLoc 2025
AgentFL 2024
CoSIL 2025
AutoFL 2024
LineVul 2022
LOVA 2024b
MatsVD 2024
xLoc 2024b
LLM4FL 2024
MemFL 2025
@@
Line Level Localization
Multiple Agents
Runtime Evidence
Iterative loop
@@ 4.1 T2L Baseline Solver
To support the proposed trace-to-line evaluation task, we provide T2L-Agent as a baseline solver that implements an end-to-end workflow from runtime traces to line-level localization under a reproducible evaluation framework. The solver uses a hierarchical planner–executor design (Udeshi et al., 2025) that splits localization into evidence gathering, hypothesis generation, and iterative refinement. Unlike single-pass analyzers, it uses a human-like process: gather runtime signals, link them to code, and narrow the search space. Figure 1 provides an overview of the end-to-end T2L workflow, including structure of T2L ARVO dataset, trace-driven evidence collection via ATA, Planner and Executor Agent, and the Evaluation running demo. Details of the tool list that T2L originally supports are provided in  A.5.
Evidence Tracing T2L Planner coordinates repository analysis and runtime evidence capture in a structured pipeline. First, code-structure analysis partitions the repositories using tree-sitter into function-aligned, semantically meaningful chunks that preserve syntactic relationships while fitting LLM context windows. Patch locations are indexed for evaluation. Next, the solver collects runtime evidence as the primary signal for trace-to-line localization task.
We collect multiple types of runtime and static evidence, including crash logs, sanitizer reports(memory-violation patterns and allocation context), stack traces, debugger snapshots, and static analyzer results. All evidence is merged into a unified textual block and fed to the LLM. This design ensures that the LLM forms its reasoning based on global evidence context, rather than aggregating per-evidence predictions. The integrated workflow is a key design in our evaluation framework and baseline solver as detailed in Agentic Trace Analyzer.
@@
Agentic Trace Analyzer (ATA) The Agentic Trace Analyzer (ATA) is the core module of T2L-Agent baseline solver. Its purpose is to connect runtime behavior with relevant source code regions by running targets in reproducible Docker, and instrumenting executions with analysis toolkit such as Sanitizers, Debuggers, and Static analyzers to collect stack traces, memory errors, and other runtime cues. The resulting evidence is constructed into a unified “evidence block” which is then passed to T2L-Agent baseline solver’s subsequent hypothesis generation and refinement steps. This ATA is designed to match the characteristics of the trace-guided, repo-scale, line-level localization task, where crash points alone are often insufficient to identify the root-cause lines.
Hypothesis Generation A “hypothesis” is a candidate vulnerable location proposed by the LLM given the unified evidence block for a repository. The LLM proposes a list of file:line candidate pairs ranked by confidence. Each hypothesis is independently checked against the ground-truth patch diff, and candidates do not affect one another during evaluation.
Feedback Control Each iteration outputs a brief task summary with success indicators and confidence. The Planner uses these signals to decide the next step: continue refining or stop, preventing premature termination and avoiding over-analysis under a fixed cost budget.
(a) Divergence Tracing
(b) Detection Refinement
@@
Figure 2: Partial T2L-Agent logs to show the how the three proposed technique on T2L-Agent work and help the task: Divergence Tracing, Detection Refinement and Agentic Trace Analyzer.